# Workload-Aware Incremental Reclustering in Cloud Data Warehouses

### v1

@@ 5.2. Implementation Details
We implement the service using DuckDB (Raasveldt, 2022) as the execution engine and deploy in a cloud environment backed by AWS S3 and Redis (39) (metadata). Data is stored in Parquet (10). During ingestion, we use Arrow (Lamb et al., 2024) to write data into fixed-size partitions and extract partition-level statistics from file headers into Redis.
@@
We consider three major optimizations in DuckDB: 1) projection pushdown, 2) column chunks, and 3) encoding and compression. For unpruned partitions, we record both the total column chunk size retrieved and the actual size read by a query in each column after projection pushdown. A partition’s utilization is then adjusted according to each column chunk’s compression ratio. These cost metrics are evaluated asynchronously to determine a reclustering decision. After a reclustering job finishes, a new table snapshot is published to keep production workloads uninterrupted.
@@ 5.3. Pruning Opportunities
Zonemap-based pruning extends beyond simple filters to support complex operations. We summarize the pruning opportunities for representative query types, as documented by Snowflake (Zimmerer et al., 2025). For LIMIT clauses, a well-clustered table facilities early termination once enough fully-matching partitions are found. For JOIN operations, build-side values are summarized into zonemaps for the probe side to prune partitions, which are essentially range queries that benefit from our reclustering approach. Snowflake builds a pruning tree to handle complex expressions across diverse predicates. In our context, each node in this tree identifies a subset of boundary partitions associated with the base columns. Our approach improves the efficiency of this pruning tree, and the resulting evaluation process guides our reclustering decisions.
@@ 6. EVALUATION
We evaluate our designs on TPC-H (Council, 2024), DSB (Ding et al., 2021a), and real-world workload Mirrors. Our principle findings are:
@@ 6.1. Workload Generator
Standard benchmarks (e.g., TPC-H) largely rely on static datasets with fixed query templates and limited dynamic features. We therefore develop a workload generator to extend standard benchmarks with a set of configurable parameters to capture the interplay between continuous data ingestion and evolving workload patterns. The generator divides a workload into consecutive periods. Each period is configured to represent a specific workload pattern and consists of a sequence of workload batches (e.g., 12 batches per period). Each batch includes a data ingestion phase, a query execution phase, and a potential reclustering phase [^5] Reclustering is invoked after each batch but may decide to take no action..
Data Ingestion. For static benchmarks such as TPC-H, we slice the dataset into chronological batches (e.g., grouped monthly by order date), preserving the original data ordering. For DSB and Mirrors (our real-world dataset), because their datasets have built-in timestamps, we use “refresh runs” in DSB and “replay logs” in Mirrors to generate the ingestion batches. The first few (with a configurable #) ingestion batches form the initial state of the database.
@@ 6.2. Workloads
We describe how the above workload generator apply to TPC-H, DSB, and Mirrors.
@@ 6.2.1. TPC-H
We generate a TPC-H dataset with a scale factor of 720. We divide the dataset into 72 monthly batches grouped by (o_orderdate) [^6] The maximum gap between l_shipdate and o_orderdate is extended from TPC-H default of 121 days to 1,000 days accordingly.. Reclustering targets the largest fact table lineitem. Each batch comprises all 8 query templates that filter on l_shipdate. The default predicate selectivity is set to a two-month interval. Each batch runs: 1) a global workload of the 8 queries with uniformly sampled predicates across the entire date range, and 2) a local workload of the 8 queries following a Zipf ($\alpha=2$) distribution skewed toward recent months.
The workload includes a total of 6 periods ($\mathbf{P1}$ to $\mathbf{P6}$), each consisting of 12 consecutive batches/months. The first two periods $\mathbf{P1,P2}$ form the initial data pool. The workload maintains a shifting rate of 25% in $\mathbf{P2}$ and $\mathbf{P3}$ ($\mathbf{P2}$ for warm up). Reclustering begins from $\mathbf{P3}$ and continues thereafter. Workload shifting rate rises to 75% from $\mathbf{P4}$. To mimic a major workload shift, the primary predicate column changes to l_commitdate in $\mathbf{P5}$, and to a 2:1 mix of l_shipdate and l_commitdate in $\mathbf{P6}$.
@@ 6.2.2. DSB
We follow the similar settings as in the TPC-H above. The workload consists of 6 periods, each containing 12 batches. An initial dataset (SF=240) and these 72 “refresh runs” ingestions bring the final scale factor to around 720. The first two periods ($\mathbf{P1,P2}$) expands the initial data pool. Reclustering begins from $\mathbf{P3}$ and targets the three largest fact tables: store_sales, catalog_sales, and web_sales. Each batch includes all 30 DSB queries that filter on these tables. Predicate values are sampled from DSB’s default Gaussian ($\sigma=2$) distribution, with dates centered on a reference month that advances by one month per batch. We set the workload shifting rate to 100% and fix query selectivity at a two-month interval.
@@ 6.2.3. Mirrors
Our real dataset, called Mirrors, comprises access logs collected from a major open-source software mirror. Spanning the last 120 days, it contains 2.4TB of raw web server logs, which compress to a total of 212GB. Each record includes source address, user agent, access path, response status code, response size, and multiple access related information fields [^7] An obfuscated example: 117.176.220.183 - - [29/Sep/2024:10:00:31 +0800] ”GET /pypi/web/packages/f6/ab/c7d5e79d2984001911d864af8ec74492da5dba558737b10774ce27587239/duckdb-1.1.0-cp310-cp310-manylinux_2_17_x86_64.manylinux2014_x86_64.whl HTTP/1.1” 200 20097722 ”application/octet-stream” ”-” ”poetry/1.8.3 CPython/3.10.15 Linux/5.15.153.1-microsoft-standard-WSL2” - https.
@@ 6.3. Experimental Setup
    No Recluster (NR): A simple baseline that keeps the data untouched in its original ingest order.
@@
    Data-Driven Incremental (DD): We choose Dremio to represent the data-driven incremental approach. Dremio uses Snowflake’s overlapping metric to rank partitions and, in each batch, reclusters up to a fixed number of partitions whose overlapping depth exceeds a predefined threshold.
@@
All experiments are conducted on the AWS cloud platform (6), using EC2 instances as compute nodes and S3 Express One Zone as the object storage. For TPC-H and DSB, we use m8gd.8xlarge instances, because smaller instance types provide unstable bandwidth. For Mirrors, we use m8gd.16xlarge instances to accommodate the increased data volume. All data transfers between EC2 and S3 are routed through S3 gateway endpoints.
@@ 6.5.1. Per-Period Breakdown
Figure 11. Centroids of Reclustered Partitions. Each point denotes a centroid identified across all evaluation periods. Point size reflects the number of partitions sharing that centroid. Color encodes the l_shipdate coordinate.
@@
DSB Breakdown.
Figure 14. Mirrors Breakdown. Left: Cumulative 10-Day Runtime. Right: CDF of Per-Query Runtime Reduction Percentages.margin:
Mirrors Breakdown.
6.5.2. Workload Characteristics
We then analyze the strengths and weaknesses of each method by comparing their different query costs under the recency-skewed local workload and the uniform global workload in TPC-H. As shown in Figure 9, DD’s gains on local is smaller than its gains on global. This is because DD treats the dataset uniformly without workload awareness. It fails to prioritize recent partitions that dominate local queries. NN achieves comparable local performance to DD but performs worse on global. Each batch is sorted in isolation, producing many disjoint “sorted runs” that fragment the table and gradually degrade clustering quality. Because NN never reclusters existing data, the initial pool remains unordered, resulting in poor performance on the global workload. QD excels on the global workload because that workload is uniformly distributed and repeats over time, making a costly one-time full repartition beneficial. However, a Qd-tree trained on historical statistics quickly becomes stale on local, which is biased toward recent data. In contrast, WAIR adapts gracefully to both workload profiles by continuously adjusting its reclustering strategy to match observed query patterns, delivering balanced and robust performance in a mixed workload.
6.5.3. Sensitivity Analysis
margin:
@@ 6.6. DSB
Despite that DSB is far more complex than TPC-H, WAIR consistently outperforms the baselines as the workload evolves, as shown in Figure 13. These results validate WAIR’s pruning effectiveness on the complex workloads. Although more non-leaf operators (e.g., join, aggregate) in a query plan add compute rather than I/O complexity directly, they are also likely to involve more predicates and columns from different base tables (and thus more pruning opportunities). WAIR’s savings signatures and hybrid layouts can help identify the most effective base column(s) for reclustering.
@@ 6.7. Mirrors
As shown in Figure 14, WAIR’s performance advantage over the baselines increases as the workload execution proceeds because WAIR continuously reclusters the most critical boundary micro-partitions adaptive to the workload. The data-driven approach (i.e., DD), on the other hand, makes suboptimal reclustering decisions in this skewed, highly selective real workload due to the lack of workload awareness. QD struggles in this scenario because a static layout optimized for yesterday’s traffic often fails to effectively prune today’s IP queries. Figure 14 (right) shows the CDF of each method’s speedup relative to the NR baseline. WAIR makes more than 60% of the queries execute at least 80% faster than the NR baseline with negligible reclustering overhead.
@@ 6.8. Selectivity Analysis
We provide a sensitivity analysis on the predicate selectivity using the TPC-H workloads. As shown in Figure 15 (left), WAIR consistently outperforms all the baselines, and the performance advantages grow as the selectivity decreases. This is because WAIR’s cost model adapts to selectivity to balance query and reclustering costs, while the baselines are unaware of selectivity. WAIR’s dominance in the Mirrors workload is partly because of the low-selectivity nature of the queries. Figure 15 (right) presents the CDF of the selectivity of Mirrors’ queries.
@@ 7. CONCLUSION
This paper advocates a clean separation between reclustering policy and clustering-key selection, classifying prior work to highlight their inflexibility. We formalize boundary micro-partitions and prove that reclustering them yields near-optimal pruning with bounded logarithmic amortized overhead. Building on this, we present the Workload-Aware Incremental Reclustering (WAIR) algorithm, which uses a sliding window of recent queries and a cloud-cost model to select high-payoff boundary partitions and reorganize them into hybrid layouts with per-group keys that maximize expected savings. margin: Across TPC-H, DSB, and a large real-world workload, WAIR cuts total cost by up to 61%, 60%, and 86%, respectively, outperforming research prototypes and documented commercial baselines. Breakdowns and sensitivity analyses show these gains are robust under dynamic conditions. In short, WAIR turns continuous, cost-effective reclustering into a practical reality for modern cloud data warehouses.

### v2

@@ 5.2. Implementation Details
We implement the service using DuckDB (Raasveldt and Mühleisen, 2019; Raasveldt, 2022) as the execution engine and deploy in a cloud environment backed by AWS S3 and Redis (40) (metadata). Data is stored in Parquet (11). During ingestion, we use Arrow (7) to write data into fixed-size partitions and extract partition-level statistics from file headers into Redis.
@@
We consider three major optimizations in DuckDB: 1) projection pushdown, 2) column chunks, and 3) encoding and compression. For unpruned partitions, we record both the total column chunk size retrieved and the actual size read by a query in each column after projection pushdown. A partition’s utilization is then adjusted according to each column chunk’s compression ratio. These cost metrics are evaluated asynchronously to determine a reclustering decision. After a reclustering job finishes, the metadata of these new partitions is recorded, and a new table snapshot is published to keep production workloads uninterrupted.
@@ 5.3. Pruning Opportunities
Zonemap-based pruning extends beyond simple filters to support complex operations. We summarize the pruning opportunities for representative query types, as documented by Snowflake (Zimmerer et al., 2025). For LIMIT clauses, a well-clustered table facilitates early termination once enough fully-matching partitions are found. For JOIN operations, build-side values are summarized into zonemaps for the probe side to prune partitions, which are essentially range queries that benefit from our reclustering approach. Snowflake builds a pruning tree to handle complex expressions across diverse predicates. In our context, each node in this tree identifies a subset of boundary partitions associated with the base columns. Our approach improves the efficiency of this pruning tree, and the resulting evaluation process guides our reclustering decisions.
@@ 6. EVALUATION
We evaluate our designs on TPC-H (Council, 2024), DSB (Ding et al., 2021a), and real-world workload Mirrors. Our principal findings are:
@@ 6.1. Workload Generator
Standard benchmarks (e.g., TPC-H) largely rely on static datasets with fixed query templates and limited dynamic features. We therefore develop a workload generator to extend standard benchmarks with a set of configurable parameters to capture the interplay between continuous data ingestion and evolving workload patterns. The generator divides a workload into consecutive periods. Each period is configured to represent a specific workload pattern and consists of a sequence of workload batches (e.g., 12 batches per period). Each batch includes a data ingestion phase, a query execution phase, and an optional reclustering phase [^5] Reclustering is invoked after each batch but may decide to take no action..
Data Ingestion. For static benchmarks such as TPC-H, we slice the dataset into chronological batches (e.g., grouped monthly by order date), preserving the original data ordering. For DSB and Mirrors (our real-world dataset), because their datasets have built-in timestamps, we use “refresh runs” in DSB and “replay logs” in Mirrors to generate the ingestion batches. An initial, configurable number of ingestion batches forms the initial state of the database.
@@ 6.2. Workloads
We describe how the above workload generator applies to TPC-H, DSB, and Mirrors.
@@ 6.2.1. TPC-H
We generate a TPC-H dataset with a scale factor of 720. We divide the dataset into 72 monthly batches grouped by (o_orderdate) [^6] The maximum gap between l_shipdate and o_orderdate is extended from TPC-H default of 121 days to 1,000 days accordingly to create a larger mismatch between ingestion order and query predicates.. Reclustering targets the largest fact table lineitem. Each batch comprises all 8 query templates that filter on l_shipdate. The default predicate selectivity is set to a two-month interval. Each batch runs: 1) a global workload of the 8 queries with uniformly sampled predicates across the entire date range, and 2) a local workload of the 8 queries following a Zipf ($\alpha=2$) distribution skewed toward recent months.
The workload includes a total of 6 periods ($\mathbf{P1}$ to $\mathbf{P6}$), each consisting of 12 consecutive batches. The first two periods $\mathbf{P1,P2}$ form the initial data pool. The workload maintains a shifting rate of 25% in $\mathbf{P2}$ and $\mathbf{P3}$ ($\mathbf{P2}$ for warm up). Reclustering begins from $\mathbf{P3}$ and continues thereafter. Workload shifting rate rises to 75% from $\mathbf{P4}$. To mimic a major workload shift, the primary predicate column changes to l_commitdate in $\mathbf{P5}$, and to a 2:1 mix of l_shipdate and l_commitdate in $\mathbf{P6}$.
@@ 6.2.2. DSB
We follow settings similar to the TPC-H setup above. The workload consists of 6 periods, each containing 12 batches. An initial dataset (SF=240) and these 72 “refresh runs” ingestions bring the final scale factor to around 720. The first two periods ($\mathbf{P1,P2}$) expand the initial data pool. Reclustering begins from $\mathbf{P3}$ and targets the three largest fact tables: store_sales, catalog_sales, and web_sales. Each batch includes all 30 DSB queries that filter on these tables. Predicate values are sampled from DSB’s default Gaussian ($\sigma=2$) distribution, with dates centered on a reference month that advances by one month per batch. We set the workload shifting rate to 100% and fix query selectivity at a two-month interval.
@@ 6.2.3. Mirrors
Our real dataset, called Mirrors, comprises access logs collected from a major open-source software mirror (46). Spanning the last 120 days, it contains 2.4TB of raw web server logs, which compress to a total of 212GB. Each record includes source address, user agent, access path, response status code, response size, and multiple access related information fields [^7] An obfuscated example: 117.176.220.183 - - [29/Sep/2024:10:00:31 +0800] ”GET /pypi/web/packages/f6/ab/c7d5e79d2984001911d864af8ec74492da5dba558737b10774ce27587239/duckdb-1.1.0-cp310-cp310-manylinux_2_17_x86_64.manylinux2014_x86_64.whl HTTP/1.1” 200 20097722 ”application/octet-stream” ”-” ”poetry/1.8.3 CPython/3.10.15 Linux/5.15.153.1-microsoft-standard-WSL2” - https (45).
@@ 6.3. Experimental Setup
    No Recluster (NR): A simple baseline that keeps the data untouched in its original ingest order without applying any reclustering.
@@
    Data-Driven Incremental (DD): We choose Dremio to represent the data-driven incremental approach. Dremio uses Snowflake’s overlapping metric to rank partitions and, in each batch, reclusters up to a fixed number of partitions whose overlapping depth exceeds a predefined overlapping threshold.
@@
All experiments are conducted on the AWS cloud platform (6), using EC2 instances as compute nodes and S3 Express One Zone as the object storage. For TPC-H and DSB, we use m8gd.8xlarge instances, because smaller instance types provide unstable bandwidth. For Mirrors, we use larger m8gd.16xlarge instances to accommodate the increased data volume. All data transfers between EC2 and S3 are routed through S3 gateway endpoints.
@@ 6.5.1. Per-Period Breakdown
Figure 12. Centroids of Reclustered Partitions. Each point denotes a centroid identified across all evaluation periods in TPC-H. Point size reflects the number of partitions sharing that centroid. Color encodes the l_shipdate coordinate value.
@@
6.5.2. Workload Characteristics
We then analyze the strengths and weaknesses of each method by comparing their different query costs under the recency-skewed local workload and the uniform global workload in TPC-H. As shown in Figure 10, DD’s gains on local are smaller than its gains on global. This is because DD treats the dataset uniformly without workload awareness. It fails to prioritize recent partitions that dominate local queries. NN achieves comparable local performance to DD but performs worse on global. Each batch is sorted in isolation, producing many disjoint “sorted runs” that fragment the table and gradually degrade clustering quality. Because NN never reclusters existing data, the initial pool remains unordered, resulting in poor performance on the global workload. QD excels on the global workload because that workload is uniformly distributed and repeats over time, making a costly one-time full repartition beneficial. However, a Qd-tree trained on historical statistics quickly becomes stale on local, which is biased toward recent data. In contrast, WAIR adapts gracefully to both workload profiles by continuously adjusting its reclustering strategy to match observed query patterns, delivering balanced and robust performance in a mixed workload.
6.5.3. Sensitivity Analysis
margin:
@@ 6.6. DSB
Although DSB is far more complex than TPC-H, WAIR consistently outperforms the baselines as the workload evolves, as shown in Figure 14. These results validate WAIR’s pruning effectiveness on the complex workloads. Although more non-leaf operators (e.g., join, aggregate) in a query plan add compute rather than I/O complexity directly, they are also likely to involve more predicates and columns from different base tables (and thus more pruning opportunities). WAIR’s savings signatures and hybrid layouts can help identify the most effective base column(s) for reclustering.
DSB Breakdown.
@@ 6.7. Mirrors
Figure 15. Mirrors Breakdown. Left: Cumulative 10-day runtime. Right: ECDF of per-query runtime reduction percentages relative to NR.margin:
Mirrors Breakdown.
@@
As shown in Figure 15, WAIR’s performance advantage over the baselines increases as the workload execution proceeds because WAIR continuously reclusters the most critical boundary micro-partitions adaptive to the workload. The data-driven approach (i.e., DD), on the other hand, makes suboptimal reclustering decisions in this skewed, highly selective real workload due to the lack of workload awareness. QD struggles in this scenario because a static layout optimized for yesterday’s traffic often fails to effectively prune today’s IP queries. Figure 15 (right) shows the ECDF of each method’s speedup relative to the NR baseline. WAIR makes more than 60% of the queries execute at least 80% faster than the NR baseline with negligible reclustering overhead.
6.8. Selectivity Analysis
@@
We provide a sensitivity analysis on the predicate selectivity using the TPC-H workloads. As shown in Figure 16 (left), WAIR consistently outperforms all the baselines, and the performance advantages grow as the selectivity decreases. This is because WAIR’s cost model adapts to selectivity to balance query and reclustering costs, while the baselines are unaware of selectivity. WAIR’s dominance in the Mirrors workload is partly because of the low-selectivity nature of the queries. Figure 16 (right) presents the ECDF of the selectivity of Mirrors’ queries.
@@ 7. CONCLUSION
This paper advocates a clean separation between reclustering policy and clustering-key selection, classifying prior work to highlight their inflexibility. We formalize boundary micro-partitions and prove that reclustering them yields near-optimal pruning with bounded logarithmic amortized overhead. Building on this, we present the WAIR algorithm and implement it into a prototype automatic reclustering service. WAIR uses a sliding window of recent queries and a cloud-cost model to select high-payoff boundary partitions and reorganize them into hybrid layouts with per-group keys that maximize expected savings. margin: Across TPC-H, DSB, and a large real-world workload, WAIR cuts total cost by up to 61%, 60%, and 86%, respectively, outperforming research prototypes and documented commercial baselines. Breakdowns and sensitivity analyses show these gains are robust under dynamic conditions. In short, WAIR turns continuous, cost-effective reclustering into a practical reality for modern cloud data warehouses.
Acknowledgements.
This paper was supported by the National Natural Science Foundation of China (Grant No. 62532001), Xiongan AI Institute, and Shanghai Qi Zhi Institute. We would also like to thank Jiaoyi Zhang, Miao Wang, and Yihao Liu for their helpful discussions and guidance.